OpenStudy (anonymous):

Which of the following represents the linear equation 2x = y + 3 in slope−intercept form?

OpenStudy (lgbasallote):

so what's your answer?

OpenStudy (anonymous):

2x=y+3 -y -y -y+2x=3 -2x -2x -y=-2x+3 divide each side by -1 y=2x-3